A spinner is divided into sections of equal size, some of which are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 25%. The probability of the arrow landing on a section colored blue is 10%. What is the probability of the arrow landing on a section colored green?
I believe it is 65% because you add 25% and 10% together then subtract from 100, because percent means per 100. So, sorry if I am wrong.
Answer:
The probability of the arrow landing on a section colored green is 65%.
Step-by-step explanation:
Given : A spinner is divided into sections of equal size, some of which are red, some are blue, and the remaining are green.
The probability of the arrow landing on a section colored red is 25%.
The probability of the arrow landing on a section colored blue is 10%.
To find : The probability of the arrow landing on a section colored green
Solution :
→ The probability of the arrow landing on a section colored red is 25%.
i.e, \(P(R)=25\%=\frac{25}{100}=0.25\)
→ The probability of the arrow landing on a section colored blue is 10%.
i.e, \(P(B)=10\%=\frac{10}{100}=0.10\)
→ Probability of the spinner is P(R)+P(B)+P(G)=1
To find the probability of section green colored, substitute the value
\(P(R)+P(B)+P(G)=1\)
\(0.25+0.10+P(G)=1\)
\(0.35+P(G)=1\)
\(P(G)=1-0.35\)
\(P(G)=0.65\)
Convert into percentage,
\(0.65=\frac{65}{100}\times 100=65\%\)
Therefore,The probability of the arrow landing on a section colored green is 65%.